# Recurrence Criteria for Reducible Homogeneous Open Quantum Walks on the Line

### v3

@@ Abstract
In this paper, we study the recurrence of Open Quantum Walks induced by finite-dimensional coins on the line ($\mathbb{Z}$) and on the grid ($\mathbb{Z}^{2}$). Two versions are considered: discrete-time open quantum walks (OQW) and continuous-time open quantum walks (CTOQW). We present three distinct recurrence criteria for OQWs on $\mathbb{Z}$, each adapted to different types of coins. The first criterion applies to coins whose auxiliary map has a unique invariant state, resulting in the first recurrence criterion for Lazy OQWs. The second one applies to Lazy OQWs of dimension 2, where we provide a complete characterization of the recurrence for this low-dimensional case. The third one presents a general criterion for finite-dimensional coins in the non-lazy case, which generalizes many of the previously known results for OQWs on $\mathbb{Z}$. Also, we present a general recurrence criterion for OQWs on $\mathbb{Z}^{2}$ via the open quantum jump chain, obtained from a recurrence criterion for CTOQWs on $\mathbb{Z}^{2}$.
@@ 1 Introduction
In this context, we first consider homogeneous Open Quantum Walks (HOQWs) defined on the set of integers, that is, the graph of this type of walk has the set of sites $V=\mathbb{Z}$, where three linear operators $L,B,R$ correspond to transitions between sites in the quantum walk, remaining constant, with each step moving only to the nearest neighbors $(R,L)$ or staying at the same site $(B)$.
Recurrence is an important probabilistic property of any walk, often characterized by the expected number of returns to the initial site, bringing a central concern in the quantum context, and can be directly related to classical recurrence criteria for walks in $\mathbb{Z}$. For example, consider the classical homogeneous walk on $\mathbb{Z}$ with graph as shown in Figure 1.
@@
In the context of discrete-time Markov chains, a state is called recurrent if, once the chain enters that state, the probability of returning to it is 1. This is equivalent to saying that the average number of returns to that state is infinite, indicating that the chain will revisit the state indefinitely [21].
With this description, it is natural to seek criteria to determine under what conditions an HOQW is recurrent, given $L,B,R.$ This problem has been addressed in the literature with the development of specific criteria, which are essential to understanding the dynamics and long-term behavior of these systems. See, for instance, [12, 17, 16]. These works define recurrence in terms of probability of return, where a state is said to be recurrent if the probability of the walk returning to it is equal to 1. However, it has been shown that for OQWs, the two definitions of recurrence based on returning with probability 1 are not equivalent to the one based on the average number of returns being infinite [4]. This property highlights a distinct characteristic of OQWs compared to classical Markov chains.
Once this difference is understood, we discuss the recurrence criteria for HOQWs, adopting the definition proposed in [4]. Specifically, we analyze the recurrence of HOQWs based on the average number of returns.
Furthermore, recurrence criteria concerning the mean number of returns for higher-order HOQWs have been presented in [15]. The authors of that work extend the analysis of recurrence to a more general class of open quantum walks, such as irreducible coins.
@@
In this article, we present more general criteria for the recurrence of lazy and non-lazy HOQWs, based on the transition operators $L,B,R$ when acting on finite-dimensional vector spaces, that is, we allow the HOQW to be reducible. Furthermore, we present the first recurrence criterion for lazy HOQWs, using results obtained from recent studies on Central Limit Theorems for this kind of HOQWs.
A first Central Limit Theorem (CLT) for Open Quantum Walks (OQWs) was established in [1]. In this work, we rely on more recent developments, including the CLT for Lazy HOQWs presented in [14], as well as the general CLT for HOQWs on lattices with finite internal degrees of freedom given in [10].
@@
By making use of the ideas above and generalizing results from the classical version to recurrence, we also obtain a recurrence criterion for the HOQW on the grid, which is the case where the set of vertices is $\mathbb{Z}^{2}$; however, this is achieved through continuous-time OQWs (CTOQWs). For this purpose, we use recent results by [20], where the introduction of the open quantum jump chain allowed for the association between recurrence in discrete-time and continuous-time versions of OQWs.
@@
This work is structured as follows. In Section 2, we define the open quantum walks induced by a coin and their transition probabilities. Section 3 is devoted to presenting a recurrence criterion for a class of Lazy HOQWs, and one restricted to Lazy HOQWs of dimension 2. In Section 4, we focus on the case where the coin is non-lazy, and we present a general recurrence criterion for finite-dimensional coins. Section 5 illustrates a first pack of comprehensive examples. Section 6 discusses OQWs on the grid, and Section 7 brings examples of HOQWs on those walks.
@@ 2 General Settings
the set of density operators acting on $\mathcal{H}\otimes\mathcal{K}.$ We are interested in HOQWs of the form
@@
We say that the triple of matrices $(L,B,R)$ is a coin if eq. (1) holds. Therefore, we can assume that $\mathfrak{h}_{i}=\mathfrak{h}\;\forall i\in\mathbb{Z},$ and if $\mathfrak{h}=\mathbb{C}^{d}$ for some $d$, the dimension of the coin is said to be $d$. In this article, we consider only the case $d<\infty.$
@@
Therefore, this coin represents a homogeneous, nearest-neighbor OQW on the integer line with the left and right transitions given by $L$ and $R$, respectively, while $B$ describes the rate of no jump at some instant.
This lazy behavior introduces distinct properties compared to its non-lazy counterpart. Notably, the presence of the matrix $B$ influences the walk’s propensity to remain in its current vertex, thereby affecting the overall dynamics. If $B=0,$ then $\Phi$ will be called a HOQW induced by a coin $(L,R),$ or a non-lazy HOQW induced by a coin $(L,R).$ The dynamic described above is represented by the graph in Figure 2.
@@
Figure 2: Lazy Quantum Walk on $\mathbb{Z}$.
@@ 3.1 Faithfulness of Density Operators
An important subset of the set of non-faithful density operators is composed of the pure density operators. A pure density is a density $\rho$ that is a projection of rank 1, i.e., there exists some vector $\lvert u\rangle\in\mathbb{C}^{d}$ such that $\rho=\lvert u\rangle\langle u\rvert$.
Let us consider an HOQW starting the walk on vertex $\lvert i\rangle$ with initial density operator $\rho.$ For evolution on time $n=0,1,2,\ldots,$ we define
@@
    $\sum_{n=0}^{\infty}p_{ii;\rho}(n)=\infty.$ (2)
@@
Concerning all vertices, an OQW is called: recurrent if every vertex is recurrent; transient if every vertex is transient. Note that equation (2) computes the mean number of returns to site $\lvert i\rangle.$
The following Proposition describes $\rho$-recurrence in a concentrated vertex $i\in\mathbb{Z}$ when assumed that $\lvert i\rangle$ is recurrent for some density and $\mathfrak{h}$ has a finite internal degree of freedom. An immediate consequence will also be given for the case where $\lvert i\rangle$ is $\sigma$-transient for some non-faithful density.
@@
is continuous and thus attains its minimum and maximum. In particular, there exists $c>0$ such that
@@
where at least one of the members of the sum must be infinity. This completes the proof.
@@
By a contra-position argument, we obtain the following direct consequence of Theorem 3.1.
@@ 3.2 Enclosures and Invariant States
Firstly, a subspace $\mathcal{Y}$ of $\mathfrak{h}$ is called an enclosure for a quantum channel $\mathcal{L}$ if for every $\sigma\in\mathcal{D}\mathfrak{h}),$ we have
@@
A minimal enclosure with finite dimension is the support of a unique invariant state of the channel, that is, a density operator $\sigma$ such that $\mathcal{L}(\sigma)=\sigma$. For more on enclosures of HOQWs, we refer the reader to [10, 11, 4].
@@
Moreover, by equation (4) we have that $\cup_{n}\;\text{supp}(\rho_{n})$ is a subspace of $\mathfrak{h}.$ Joining all the possible initial states in which $(L,B,R)$ is transient, we obtain a decomposition (3), where $\mathcal{Y}$ or $\mathcal{X}$ may be $\{0\}$, and by definition $\mathcal{Y}$ is an enclosure.
@@
In our subsequent discussion, we consider a coin $(L,B,R)$ and its auxiliary map $\mathcal{L}:\mathcal{B}(\mathfrak{h})\rightarrow\mathcal{B}(\mathfrak{h})$ given by
We will use the following result obtained straightforwardly from [14].

### v4

@@ Abstract
In this paper, we study the recurrence of Open Quantum Walks induced by finite-dimensional coins on the line ($\mathbb{Z}$) and on the grid ($\mathbb{Z}^{2}$). Two versions are considered: discrete-time open quantum walks (OQW) and continuous-time open quantum walks (CTOQW). We present three distinct recurrence criteria for OQWs on $\mathbb{Z}$, each adapted to different types of coins. The first criterion applies to coins whose auxiliary map has a unique invariant state, resulting in the first recurrence criterion for Lazy OQWs. The second applies to Lazy OQWs of dimension 2, where we provide a complete characterization of the recurrence for this low-dimensional case. The third one presents a general criterion for finite-dimensional coins in the non-lazy case, which generalizes several of the previously known results for OQWs on $\mathbb{Z}$. Also, we present a general recurrence criterion for OQWs on $\mathbb{Z}^{2}$ via the open quantum jump chain, obtained from a recurrence criterion for CTOQWs on $\mathbb{Z}^{2}$.
@@ 1 Introduction
OQWs have proved to be a versatile framework for describing discrete-time quantum dynamics driven by environmental interaction. Beyond their structural simplicity, they provide an effective language for dissipative implementations of logical operations and state engineering. In particular, it has been shown that single-qubit gates and the CNOT gate can be realized as OQWs on fully connected graphs, and that arbitrary single-qubit states as well as Bell states can be prepared through dissipative mechanisms [3]. From a probabilistic viewpoint, central limit theorems (CLTs) have been established for microscopically derived Lazy OQWs [16], as well as for the CTOQWs [8].
At the same time, OQWs admit a rich potential-theoretic and asymptotic analysis. Fundamental quantities such as passage times, expected number of visits, recurrence and transience have been investigated [4]. Concerning the implementation of dissipative quantum computing protocols based on the OQW formalism, the approach [26] enables faster convergence to the desired steady state and increases the probability of detecting the computational outcome compared to the canonical approach.
More recently a quantum PageRank algorithm based on discrete-time OQWs with Weyl operators as Kraus operators was introduced, demonstrating faster convergence properties than several existing proposals and extending ranking methods to complex networks [14]. Recurrence properties of CTOQWs were investigated in [22], where connections with discrete-time OQWs were established and recurrence criteria were derived. Complementarily, [21] analyzed CTOQWs in one dimension through a matrix framework, characterizing site recurrence and related statistics via matrix-valued orthogonal polynomials and Lindblad generators modeling quantum birth-death processes. Finally, [24] examined hitting-time statistics under quantum channels, providing practical tools to compute first-visit expectations and variances through monitored quantum evolutions and induced absorbing quantum Markov chains.
In this work, we first consider homogeneous Open Quantum Walks (HOQWs) defined on the set of integers, that is, the graph of this type of walk has the set of sites $V=\mathbb{Z}$, where three linear operators $L,B,R$ correspond to transitions between sites in the quantum walk, remaining constant, with each step moving only to the nearest neighbors $(R,L)$ or staying at the same site $(B)$.
Recurrence is an important probabilistic property of any walk, often characterized by the expected number of returns to the initial site. This is a central concern in the quantum context, and can be directly related to classical recurrence criteria for walks on $\mathbb{Z}$. For example, consider the classical homogeneous walk on $\mathbb{Z}$ with graph as shown in Figure 1.
@@
In the context of discrete-time Markov chains, a state is called recurrent if, once the chain enters that state, the probability of returning to it is 1. This is equivalent to saying that the average number of returns to that state is infinite, indicating that the chain will revisit the state indefinitely [25]. The definition of recurrence employed on this work follows the terminology of [4]. As shown therein, this notion is better suited to the quantum setting than the classical “probability of return equals 1” criterion, since for OQWs the two are not equivalent. Moreover, for irreducible semifinite OQWs, the average number of returns satisfies a universality property: it is either finite for all vertices and all initial internal states, or infinite for all vertices and all initial internal states. The same does not hold for the probability of return, which may equal 1 for some internal states and be strictly less than 1 for others, even when the walk starts at the same site. The mean return definition also connects naturally to the limit theorems used throughout this work and allows recurrence to be characterized in terms of a drift parameter $m$.
With this description, it is natural to seek criteria to determine under what conditions an HOQW is recurrent. This problem has been addressed in the literature with the development of specific criteria, which are essential to understanding the dynamics and long-term behavior of these systems. See, for instance, [13, 18, 19]. In this context, recurrence criteria for higher-order HOQWs have been established in [17], where the authors derive such criteria for non-lazy HOQWs with two-dimensional coins as well as for arbitrary finite dimensions in the case of irreducible coins.
@@
In this article, we present more general criteria for the recurrence of lazy and non-Lazy HOQWs, based on the transition operators $L,B,R$ when acting on finite-dimensional vector spaces, that is, we allow the HOQW to be reducible. Furthermore, we present the first recurrence criterion for Lazy HOQWs, using results obtained from recent studies on CLTs for this kind of HOQWs. A first CLT for OQWs was established in [1]. In this work, we rely on more recent developments, including the CLT for Lazy HOQWs presented in [16], as well as the general CLT for HOQWs on lattices with finite internal degrees of freedom given in [11].
@@
By making use of the ideas above and generalizing results from the classical version to recurrence, we also obtain a recurrence criterion for the HOQW on the grid, which is the case where the set of vertices is $\mathbb{Z}^{2}$; however, this is achieved through continuous-time OQWs (CTOQWs). For this purpose, we use recent results of [22], where the introduction of the open quantum jump chain allowed for the association between recurrence in discrete-time and continuous-time versions of OQWs.
@@
This work is structured as follows. In Section 2, we define the OQWs induced by a coin and their transition probabilities. Section 3 is devoted to presenting a recurrence criterion for Lazy HOQWs under ergodicity. Section 4 discusses recurrence under ergodicity for discrete-time and continuous-time OQWs on the grid. General recurrence criteria for HOQWs on the line and on the grid are detailed on Section 5. Section 6 illustrates a pack of comprehensive examples.
While a unified criterion encompassing both laziness and reducibility is theoretically desirable, the primary goal of this paper is to elucidate the recurrence properties of lazy walks in the ergodic regime. The effect of reducibility is addressed exclusively in the non-lazy regime. A combined treatment would require a more involved technical framework that could obscure these distinct contributions. For this reason, we have chosen to present them separately, leaving the general lazy reducible case as a natural direction for future research.
For the reader’s convenience, we summarize in Table 1 the main recurrence criteria for discrete-time HOQWs established throughout this paper, indicating for each result the graph ($\mathbb{Z}$ or $\mathbb{Z}^{2}$), the dimension of the internal space, whether the walk is lazy or non-lazy, and whether it applies to the ergodic or general (reducible) case. For continuous-time HOQWs on the grid, the main result is Theorem 4.13, which provides a recurrence criterion under the assumption that the auxiliary map has at most one minimal enclosure.
Criterion Graph Dimension Lazy/Non-Lazy Type
Theorem 3.8 $\mathbb{Z}$ Finite $d$ Lazy Ergodic
Theorem 5.2 $\mathbb{Z}$ $d=2$ Lazy General
Theorem 5.6 $\mathbb{Z}$ Finite $d$ Non-Lazy General
Theorem 4.15 $\mathbb{Z}^{2}$ Finite $d$ Non-Lazy Ergodic
Theorem 5.9 $\mathbb{Z}^{2}$ Finite $d$ Non-Lazy General
Table 1: Summary of discrete-time recurrence criteria.
@@ 2.1 Lazy Open Quantum Walks
the set of density operators acting on $\mathcal{H}\otimes\mathcal{K}.$ Also, given a vector space $T$, we denote the set of bounded operators acting on $T$ by $\mathcal{B}(T)$.
We are interested in HOQWs of the form
@@
We say that the triple of matrices $(L,B,R)$ is a coin if eq. (1) holds. Therefore, we can assume that $\mathfrak{h}_{i}=\mathfrak{h}\;\forall i\in\mathbb{Z},$ and if $\mathfrak{h}=\mathbb{C}^{d}$ for some $d$, the dimension of the coin is said to be $d$. In this article, we consider only the finite-dimensional case $d<\infty.$
@@
Therefore, this coin represents a homogeneous, nearest-neighbor OQW on the integer line with the left and right transitions given by $L$ and $R$, respectively, while $B$ describes the rate of no jump at some instant. This lazy behavior introduces distinct properties compared to its non-lazy counterpart. Notably, the presence of the matrix $B$ influences the walk’s propensity to remain in its current vertex, thereby affecting the overall dynamics. If $B=0,$ then $\Phi$ will be called a HOQW induced by a coin $(L,R),$ or a non-Lazy HOQW induced by a coin $(L,R).$ The dynamic described above is represented by the graph in Figure 2.
@@
Figure 2: Lazy HOQW on $\mathbb{Z}$.
@@
While irreducibility has been the focus of previous studies on the asymptotic behavior of HOQWs, this work extends the analysis to the reducible case. We now recall the distinction between these two notions.
We consider a coin $(L,B,R)$ and its auxiliary map $\mathcal{L}:\mathcal{B}(\mathfrak{h})\rightarrow\mathcal{B}(\mathfrak{h})$ given by
As in classical Markov chains, an irreducible HOQW cannot be decomposed into dynamically independent subsystems, ensuring well-defined ergodic properties [12].
Proposition 2.1.
[12, Proposition 4.1] Let $\Phi$ be an HOQW defined by transition operators $L_{s}$, and $\mathcal{L}$ its auxiliary map.
  1. 1.
    $\mathcal{L}$ is irreducible if and only if the operators $\{L_{s},s\in S\}$ have no nontrivial closed invariant subspace in common.
  2. 2.
    $\Phi$ is irreducible if and only if all possible concatenations of transitions starting and ending at the same site have no nontrivial closed invariant subspace in common.
A direct consequence is the following:
Corollary 2.2.
[12, Corollary 4.2] If $\Phi$ is irreducible, then $\mathcal{L}$ is irreducible.
This corollary implies that if the auxiliary map $\mathcal{L}$ is reducible, then the HOQW $\Phi$ is also reducible. As noted in [12, Remark 3.6], an irreducible map admits at most one invariant state, and if such a state exists it must be faithful. Conversely, if a map has a unique faithful invariant state, then it is irreducible. This connection between algebraic structure and spectral properties provides a practical way to verify irreducibility in concrete examples.
For instance, let $\mathfrak{h}=\mathbb{C}^{2}$, $S=\{-1,+1\}$, and denote $L_{-1}=L$, $L_{+1}=R,$ where
$L=\frac{1}{\sqrt{20}}\begin{bmatrix}4&-1\\ 0&2\end{bmatrix},\qquad R=\frac{1}{\sqrt{20}}\begin{bmatrix}2&2\\ 0&\sqrt{11}\end{bmatrix}$.
The auxiliary map $\mathcal{L}$ has a unique invariant state $\rho_{\infty}=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}$, which is not faithful. Therefore, $\mathcal{L}$ is reducible, and consequently $\Phi$ is also reducible.
@@ 3.1 Faithfulness of Density Operators
An important subset of non-faithful density operators is composed of the pure density operators. A pure density is a density $\rho$ that is a projection of rank 1, i.e., there exists some vector $\lvert u\rangle\in\mathbb{C}^{d}$ such that $\rho=\lvert u\rangle\langle u\rvert$.
Let us consider an HOQW starting the walk at vertex $\lvert i\rangle$ with initial density operator $\rho.$ For evolution at time $n=0,1,2,\ldots,$ we define
@@
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{E}_{i,\rho}(n_{i}):=}\sum_{n=0}^{\infty}p_{ii;\rho}(n)=\infty,\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{ where }p_{ii;\rho}(0)=1.}$ (2)
@@
Concerning all vertices, an OQW is called: recurrent if every vertex is recurrent; transient if every vertex is transient. Note that equation (2) computes the mean number of returns to site $\lvert i\rangle.$ The choice of this definition, based on the mean number of returns, rather than the probability of return equals 11 ($\mathbb{P}_{i,\rho}(t_{i}<\infty)<1$, where $t_{i}$ is the returning time to $i$) definition, requires a careful justification. As demonstrated in [4], for OQWs these two notions are not equivalent, unlike the classical case. This non-equivalence makes a conceptual choice necessary regarding which definition to adopt.
The definition adopted in this work, based on the mean number of returns $\mathbb{E}_{i,\rho}(n_{i})$, follows the terminology proposed in [4] for semifinite irreducible OQWs. As established in [4, Corollary 3.10], for irreducible semifinite OQWs the value of $\mathbb{E}_{i,\rho}(n_{i})$ is universal: it is either finite for every $i\in V$ and every $\rho\in\mathcal{D}(\mathfrak{h}_{i})$, or infinite for every $i$ and every $\rho$. This universality does not hold for the return probability $\mathbb{P}_{i,\rho}(t_{i}<\infty)$, which may assume different values for different internal states even in irreducible OQWs (as illustrated in [4, Example 5.2]). Therefore, the definition based on the mean number of visits is more suitable for a robust and consistent classification of asymptotic behavior, connecting naturally to the limit theorems used throughout this work and allowing the characterization of recurrence via the drift parameter $m$, as developed in the following sections.
It is worth recalling that in the particular case of irreducible HOQWs, Jacq and Lardizabal [17] demonstrated that these two notions of recurrence do coincide. However, their proof relies heavily on irreducibility assumptions, particularly on the uniqueness and faithfulness of the invariant state, which forces the equivalence between the divergence of the mean number of returns and the certainty of return. For reducible HOQWs, the relationship between these two definitions remains an open problem. The existence of multiple invariant states and the possibility of nontrivial transient components within the internal space $\mathfrak{h}$ may lead to situations where a vertex is $\rho$-recurrent under one definition but not the other, depending on the initial internal state. This question, while beyond the scope of the present work, represents an interesting direction for future investigation.
Having clarified the choice of recurrence definition and its nuances, we now proceed to the following proposition, which describes $\rho$-recurrence in a concentrated vertex $i\in\mathbb{Z}$ when assumed that $\lvert i\rangle$ is recurrent for some density and $\mathfrak{h}$ has a finite internal degree of freedom. This result establishes the behavior of $\rho$-recurrence when recurrence occurs for at least one initial state. In the following sections, we will build upon this to analyze the structure of the internal space $\mathfrak{h}$ and understand how it decomposes in order to develop recurrence criteria for reducible OQWs. An immediate consequence will also be given for the case where $\lvert i\rangle$ is $\sigma$-transient for some non-faithful density.
@@
is upper semicontinuous and thus attains its minimum and maximum. In particular, there exists $c>0$ such that
@@
where at least one term in the sum must be infinite. This completes the proof.
@@
By a contraposition argument, we obtain the following direct consequence of Theorem 3.1.
@@ 3.2 Enclosures and Invariant States
A subspace $\mathcal{Y}$ of $\mathfrak{h}$ is called an enclosure for a quantum channel $\mathcal{L}$ if for every $\sigma\in\mathcal{D}(\mathfrak{h}),$ we have
@@
A minimal enclosure with finite dimension is the support of a unique invariant state of the channel, that is, a density operator $\sigma$ such that $\mathcal{L}(\sigma)=\sigma$ [10, Section 2.2]. For more on enclosures of HOQWs, we refer the reader to [10, 11, 4, 15].
@@
[…]
@@
3.3 Recurrence under Ergodicity
Following the CLT approach for lazy OQWs given in [16], we consider the Markov chain $(\rho_{n},\Delta X_{n})_{n\geq 0}$ associated with the quantum trajectory description of the walk. Let $m\in\mathbb{R}^{d}$ denote the drift vector of the walk, given by the expected value of the position increment under the invariant state $\rho_{\infty}$, namely
$m=\sum_{k}\mathrm{Tr}(A_{k}\rho_{\infty}A_{k}^{*})\,e_{k},$
where $(e_{k})_{k}$ are the possible displacement vectors of the walk, $A_{k}$ are the transition operators. For a fixed vector $l\in\mathbb{R}^{d}$, the quantity $(X_{n}-nm)\cdot l$ admits a Doob decomposition into the sum of a martingale part and a predictable process. The martingale term is given by
$M_{n}=\sum_{j=2}^{n}\left[f(\rho_{j},\Delta X_{j})-Pf(\rho_{j-1},\Delta X_{j-1})\right],$ (5)
where $f:\mathcal{D}(\mathfrak{h})\times\mathbb{R}^{d}\to\mathbb{R}$ is defined by
$f(\rho,x)=\mathrm{Tr}(\rho L_{l})+x\cdot l,$
and $L_{l}=L\cdot l$ denotes the operator associated with the vector $l\in\mathbb{R}^{d}$ acting on the internal Hilbert space $\mathfrak{h}$. Here $\rho_{j}$ denotes the internal state of the walker at time $j$, $\Delta X_{j}=X_{j}-X_{j-1}$ represents the increment of the position process, and $P$ is the transition operator of the Markov chain $(\rho_{n},\Delta X_{n})$.